if a body move with u velocity then after collide with wall and remove with v velocity what is the impulse

Grade:12th pass

1 Answers

Ankit Jaiswal
7 years ago
Hi aman,
             I think this might help
as we know that Impulse
J = \int F.dt   
and also F=m(dv/dt) so                                                             (here m is the mass of body and dv/dt =acc)  J = \int m(\frac{dv}{dt}).dt
\therefore J = \int m.dv
\therefore J = \Delta \overrightarrow{p}
here p is the momentum ans delta p is change in momentum
hence in the above question using
impule = change in momentum
           = m\overrightarrow{v}-m\overrightarrow{u} =m(\overrightarrow{v}-\overrightarrow{u})
 
plz approve if u found it usefull 
thanks